# Deep Learning of Multivariate Extremes via a Geometric Representation

### v1

@@ 2.2 Geometric extremes on Laplace margins
As noted in Section 1, we consider data on standard Laplace margins as this permits a more detailed description of joint tail behaviour. However, the theoretical results provided by Nolde and Wadsworth, (2022), linking the unit-level set $\partial\mathcal{G}$ to a variety of modelling frameworks, are given only for random vectors on standard exponential margins. The next proposition illustrates that some of the same results also hold for data on Laplace margins.
@@ 2.3 Constructing valid unit-level sets
We now consider estimation of the unit-level set $\partial\mathcal{G}$ or, equivalently, the gauge function $g(\cdot)$ on $\mathcal{S}^{d-1}$. This requires estimates of the corresponding limit set, $\mathcal{G}$, to have certain properties; see Section 2.1. In what follows, we show that such properties are easily satisfied via an appropriate estimator for $g(\cdot)$. We begin by noting that $g(\cdot)$ must satisfy the constraint described in the following Proposition 2.2.
@@
Ignoring this constraint may lead to estimates of $\mathcal{G}$ not contained within $[-1,1]^{d}$. Such estimates lack any theoretical interpretation and must undergo a user-specified rescaling to be of use; see, e.g., Papastathopoulos et al., (2024) or Murphy-Barltrop et al., 2024a . We exploit Proposition 2.2 to ensure that our model does not suffer from this problem and can be designed to always provide valid estimates of limit sets. Now consider any continuous radial function $h:\mathcal{S}^{d-1}\mapsto\mathbb{R}_{+}$. The following proposition holds.
@@
Proof of Proposition 2.3 is provided in Appendix B.2.
Proposition 2.3 implies that we can design models for $\mathcal{G}$ that satisfy (some of) the validity properties of limit sets by starting with any (potentially arbitrary) continuous radial function $h(\cdot)$ that satisfies $1/h(\bm{w})\geq||\bm{w}||_{\infty}$; the corresponding set $\mathcal{H}$ will be a subset of $[-1,1]^{d}$, star-shaped, and compact. Whilst $\mathcal{H}$ must be a subset of the unit hyper-cube, it is not guaranteed to intersect with its boundary in each component, or even at all; the same is true of the boundary set $\partial\mathcal{H}$. Since this is a requirement of valid limit and unit-level sets, we propose a rescaling procedure to ensure it is always satisfied for our model.
@@
The rescaled set $\widetilde{\partial\mathcal{H}}$ is in one-to-one correspondence with $\partial\mathcal{H}$, satisfies ~$\widetilde{\partial\mathcal{H}}\subset[-1,1]^{d}$, and has componentwise maxima and minima $\bm{1}_{d}$ and $-\bm{1}_{d}$, respectively.
@@
Lemma 2.1 is used to prove Proposition 2.4 in Appendix B.3. Proposition 2.4 permits a new construction of valid gauge functions, denoted by $\tilde{g}:\mathcal{S}^{d-1}\mapsto\mathbb{R}_{+}$, with
@@
The rescaled gauge function $\tilde{g}(\cdot)$ satisfies $\tilde{g}(\bm{w})\geq||\bm{w}||_{\infty}$ for all $\bm{w}\in\mathcal{S}^{d-1}$. Let $\bm{w}^{u,i}=\operatorname*{arg\,max}_{\bm{w}\in\mathcal{S}^{d-1}}\{w_{i}h(\bm{w})\}$ and $\bm{w}^{l,i}=\operatorname*{arg\,min}_{\bm{w}\in\mathcal{S}^{d-1}}\{w_{i}h(\bm{w})\}$ for $i=1,\dots,d$. Then $\tilde{g}(\kappa(\bm{w}^{u,i}))=||\kappa(\bm{w}^{u,i})||_{\infty}$ and $\tilde{g}(\kappa(\bm{w}^{l,i}))=||\kappa(\bm{w}^{l,i})||_{\infty}$ for each $i=1,\ldots,d$.
@@
For inference of the limit set, we exploit Corollary 2.2 to ensure our estimates of the limit (or unit-level) set have the required properties for validity; see Section 3.2. We note that an alternative rescaling was proposed by Papastathopoulos et al., (2024). However, this was applied post-hoc on an initial estimate of the gauge function via a two-step procedure, whereas our rescaling is performed during inference without the need for an additional step.
@@ 2.4 Extended Angular Dependence Function
Consider now any gauge function $g(\cdot)$ with limit and unit level sets $\mathcal{G}$ and $\partial\mathcal{G}$ respectively. Proposition 2.1 implies that valid estimates of $\partial\mathcal{G}$ can be used to immediately obtain parameter estimates for several existing modelling frameworks, providing information about the extremal dependence structure in the positive orthant of $\mathbb{R}^{d}$. However, as noted in Section 1, our interest lies more generally in understanding the extremal dependence in all orthants of $\mathbb{R}^{d}$. The next proposition shows that, with suitable scaling coefficients, the limit set in any orthant can be obtained by considering a rescaled random vector on the set $\mathcal{S}^{d-1}_{+}$.
@@
where $\mathcal{L}(\cdot;\bm{w})$ is a slowly varying function and $\Lambda(\cdot)$ is defined in Proposition 2.8. Note that the extended ADF corresponds with the copula exponent function introduced by Mackay and Jonathan, (2023) for data on uniform margins. The next proposition illustrates that, under mild conditions, the convergence in (6) is always achieved.
@@
Proof of Proposition 2.7 is provided in Appendix B.6. Note that while the assumptions for Proposition 2.7 may seem restrictive, Wadsworth and Tawn, (2013) demonstrate through rigorous theoretical treatment that model (2) captures the joint tail structure of $\bm{X}_{E}$ for a wide variety of theoretical examples. Therefore, it is reasonable to assume this framework can capture the tail structure in every orthant.
@@
In the definition of our model (6), we purposely exclude any angles in $\mathcal{S}^{d-1}$ that intersect the axes, as the model is not well defined there. Take, e.g., $\bm{w}:=(1,0,\ldots,0)^{T}$; it is not clear whether one should consider the probability $\Pr(X_{1}>u,X_{i}>0,i=2,\ldots,d)$ or $\Pr(X_{1}>u,X_{i}<0,i=2,\ldots,d)$ under the modelling framework. This results in discontinuities for the extended ADF at the axes; see Figure 2.
@@
Proof of Corollary 2.3 is provided in Appendix B.8, with a visualisation provided in Figure 2. The blue lines in each panel of Figure 2 denote the sets $\{\bm{w}/\Lambda(\bm{w}):\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}\}$, corresponding to the intersection points $\{\tilde{\mathfrak{r}}_{\bm{w}}\tilde{\mathcal{R}}_{\bm{w}}\cap\partial\mathcal{G},\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}$}. Furthermore, from the first and third panels of Figure 2, one can clearly observe the discontinuities of the extended ADF at the axes, as discussed in Remark 4.
@@ 3.2 Modelling assumptions for the conditional radii
Wadsworth and Campbell, (2024) demonstrate that for large radial values $r$, we have that $f_{R\mid\bm{W}}(r\mid\bm{w})\propto r^{d-1}\exp\{-rg(\bm{w})\}$, where $f_{R\mid\bm{W}}$ denotes the density function of $R\mid\bm{W}$. This implies that the upper tail of $R\mid(\bm{W}=\bm{w})$ should follow the form of a gamma kernel. Note that this holds for $\bm{X}$ on both exponential and Laplace margins. To accommodate this form, Wadsworth and Campbell, (2024) propose the modelling assumption:
@@
where ‘truncGamma’ denotes a truncated Gamma distribution with shape and rate parameters $\alpha>0$ and $g(\bm{w})>0$, respectively, and $r_{\tau}(\bm{w})$ denotes the $\tau$-quantile of $R\mid(\bm{W}=\bm{w})$ for some $\tau\in(0,1)$ close to 1, that is, $\Pr\{R\leq r_{\tau}(\bm{w})\mid\bm{W}=\bm{w}\}=\tau$ for all $\bm{w}\in\mathcal{S}^{d-1}$. Using form (7), we can use maximum likelihood techniques to estimate the gauge function $g(\cdot)$, which is viewed as a non-stationary rate parameter of the truncated gamma distribution.
@@ 3.3 Estimation of the gauge function using neural networks
To overcome the limited flexibility of existing geometric modelling approaches, we model the rescaled gauge function ($\tilde{g}$; see Corollary 2.1) using neural networks. Full inference requires the construction of two models: one for the radial threshold, $r_{\tau}(\bm{w})$ in (7), and one for an (unscaled) gauge function ${g}(\bm{w})$ satisfying ${g}(\bm{w})\geq||\bm{w}||_{\infty}$ for all $\bm{w}\in\mathcal{S}^{d-1}$ (see Proposition 2.2). With the latter, we can employ the transformation in Corollary 2.1 to produce a model for the rescaled gauge function, $\tilde{g}(\bm{w})$, that provides a valid unit-level set.
@@
where we have suppressed the dependency of $r_{\tau}(\cdot)$ on $\bm{\psi}$. As $r_{\tau}(\bm{w})$ denotes the $\tau$-quantile of $R\mid(\bm{W}=\bm{w})$, the most appropriate choice of loss function is the tilted loss, given by $l(r,r_{\tau})=\rho_{\tau}(r-r_{\tau})$ for $\rho_{\tau}(z):=z(\tau-\mathbbm{1}\{z<0\})$ (Koenker et al.,, 2017). We can also define a suitable loss function and minimisation problem to train the deep representation of the rescaled gauge function, $\tilde{g}(\bm{w}).$ This model can be described as a conditional density network estimation (see, e.g., Rothfuss et al.,, 2019), where the loss function is the negative log-likelihood associated with the truncated gamma model defined in (7); note that this is dependent on both the exceedance threshold $r_{\tau}(\bm{w}_{j})$ and a scaling parameter, $\alpha>0$. To fit the gauge function model, we replace $\ell(\cdot,\cdot)$ in (8) with
@@
Full inference for our framework is performed in a two-stage fashion. We first train a model for $r_{\tau}(\bm{w})$ and derive its estimate, $\hat{r}_{\tau}(\bm{w})$. This estimate is then using in a subsequent training step for the rescaled gauge function $\tilde{g}(\cdot)$ and $\alpha$, by replacing $r_{\tau}(\bm{w})$ in (3.3) with its estimated counterpart. For details of the algorithms and regularisation techniques used for training of both models, as well as practical advice for pre-training (Goodfellow et al.,, 2016) of the models using initial estimates for $r_{\tau}(\cdot)$, see Appendix A.2.
@@ 3.4 Estimating the extended ADF
Estimates $\hat{\Lambda}(\bm{w}),\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A},$ can be used to obtain probability estimates for a wide variety of joint tail regions of $\bm{X}$. The estimation scheme outlined below avoids the need to sample from, or model, the distribution of the angles $\bm{W}$, as has been considered in other probability estimation schemes using the geometric representation (Papastathopoulos et al.,, 2024; Wadsworth and Campbell,, 2024). This feature is practically advantageous as estimation of the angular distribution alongside the limit set introduces additional modelling uncertainty into our framework, and many existing techniques for non-parametric density estimation are only applicable in low dimensional settings (Ruzgas et al.,, 2021).
To begin, let $\bm{x}\in\mathbb{R}^{d}$ be such that $r:=||\bm{x}||$ is large and define the corresponding angle $\bm{w}:=\bm{x}/r$. Further define the univariate structure variable $T_{\bm{w}}:=\min_{i=1,\dots,d}\{X_{i}/w_{i}\}$, and let $u$ denote a quantile of $T_{\bm{w}}$ satisfying $q:=\Pr(T_{\bm{w}}\leq u)<\Pr(T_{\bm{w}}\leq r)$, with $q$ close to 1 and $u>0$ large. Equation (6) implies that
@@
$=\Pr(T_{\bm{w}}>r\mid T_{\bm{w}}>u)\Pr(T_{\bm{w}}>u)\approx\exp\{-\hat{\Lambda}(\bm{w})(r-u)\}(1-q).$ (11)
Figure 4 illustrates examples of joint tail regions in the two dimensional setting that can be estimated using the framework described in equation (6). Observe that these regions are more general than many inference procedures for multivariate extremes, which just focus on the joint survivor or cumulative distribution function (e.g., Ledford and Tawn,, 1996; Ramos and Ledford,, 2009; Cooley et al.,, 2019).

### v2

@@ 2.2 Geometric extremes on Laplace margins
As noted in Section 1, we consider data on standard Laplace margins as this permits a more detailed description of joint tail behaviour. However, the theoretical results provided by Nolde and Wadsworth, (2022) linking the unit-level set $\partial\mathcal{G}$ to a variety of modelling frameworks are given only for random vectors on standard exponential margins. The next proposition illustrates that some of the same results also hold for data on Laplace margins.
@@ 2.3 Constructing valid unit-level sets
We now consider estimation of the unit-level set $\partial\mathcal{G}$ or, equivalently, the gauge function $g(\cdot)$ on $\mathcal{S}^{d-1}$. This requires estimates of the corresponding limit set, $\mathcal{G}$, to have certain properties; see Section 2.1. In what follows, we show that such properties are easily satisfied via an appropriate estimator for $g(\cdot)$.
We begin by noting that $g(\cdot)$ must satisfy the constraint described in the following Proposition 2.2.
@@
Ignoring this constraint may lead to estimates of $\mathcal{G}$ not contained within the hypercube $[-1,1]^{d}$. Such estimates lack any theoretical interpretation and must undergo a user-specified rescaling to be of use; see, e.g., Papastathopoulos et al., (2024) or Murphy-Barltrop et al., 2024a . We exploit Proposition 2.2 to ensure that our model does not suffer from this problem and can be designed to always provide valid estimates of limit sets. Now consider any continuous radial function $h:\mathcal{S}^{d-1}\mapsto\mathbb{R}_{+}$. The following proposition holds.
@@
Proof of Proposition 2.3 is provided in Appendix B.2. Proposition 2.3 implies that we can design models for $\mathcal{G}$ that satisfy (some of) the validity properties of limit sets by starting with any (potentially arbitrary) continuous radial function $h(\cdot)$ that satisfies $1/h(\bm{w})\geq||\bm{w}||_{\infty}$; the corresponding set $\mathcal{H}$ will be a subset of $[-1,1]^{d}$, star-shaped, and compact. Whilst $\mathcal{H}$ must be a subset of the unit hyper-cube, it is not guaranteed to intersect with its boundary in each component, or even at all; the same is true of the boundary set $\partial\mathcal{H}$. Since this is a requirement of valid limit and unit-level sets, we propose a rescaling procedure to ensure it is always satisfied for our model.
@@
The rescaled set $\widetilde{\partial\mathcal{H}}$ is in one-to-one correspondence with $\partial\mathcal{H}$, satisfies ~$\widetilde{\partial\mathcal{H}}\subset[-1,1]^{d}$, and has componentwise maxima and minima of $\bm{1}_{d}$ and $-\bm{1}_{d}$, respectively.
@@
Lemma 2.1 is used to prove Proposition 2.4 in Appendix B.3. Proposition 2.4 permits a new construction of a valid gauge function, denoted by $\tilde{g}:\mathcal{S}^{d-1}\mapsto\mathbb{R}_{+}$, with
@@
The rescaled gauge function $\tilde{g}(\cdot)$ satisfies $\tilde{g}(\bm{w})\geq||\bm{w}||_{\infty}$ for all $\bm{w}\in\mathcal{S}^{d-1}$. Furthermore, letting $\bm{w}^{u,i}=\operatorname*{arg\,max}_{\bm{w}\in\mathcal{S}^{d-1}}\{w_{i}h(\bm{w})\}$ and $\bm{w}^{l,i}=\operatorname*{arg\,min}_{\bm{w}\in\mathcal{S}^{d-1}}\{w_{i}h(\bm{w})\}$ for $i=1,\dots,d$, we have $\tilde{g}(\kappa(\bm{w}^{u,i}))=||\kappa(\bm{w}^{u,i})||_{\infty}$ and $\tilde{g}(\kappa(\bm{w}^{l,i}))=||\kappa(\bm{w}^{l,i})||_{\infty}$ for each $i=1,\ldots,d$.
@@
For inference of the limit set, we exploit Corollary 2.2 to ensure our estimates of the limit (or unit-level) set have the required properties for validity; see Section 3.2. We note that an alternative rescaling was proposed by Papastathopoulos et al., (2024). However, this was applied post-hoc on an initial estimate of the gauge function via a two-step procedure, whereas our rescaling is performed during inference without the need for an additional step. Furthermore, incorporating the rescaling described in equation (5) within many standard optimisation algorithms would likely result in slow model fitting procedures due to additional labor-heavy computations. However, the use of stochastic gradient descent (see Section A.2) within deep learning procedures avoids the bulk of this computational burden, allowing rescaling to be incorporated within the proposed DeepGauge framework.
@@ 2.4 Extended Angular Dependence Function
Consider now any gauge function $g(\cdot)$ with limit and unit-level sets $\mathcal{G}$ and $\partial\mathcal{G}$ respectively. Proposition 2.1 implies that valid estimates of $\partial\mathcal{G}$ can be used to immediately obtain parameter estimates for several existing modelling frameworks, providing information about the extremal dependence structure in the positive orthant of $\mathbb{R}^{d}$. However, as noted in Section 1, our interest lies more generally in understanding the extremal dependence in all orthants of $\mathbb{R}^{d}$. The next proposition shows that, with suitable scaling coefficients, the limit set in any orthant can be obtained by considering a rescaled random vector on $\mathcal{S}^{d-1}_{+}$.
@@
where $\mathcal{L}(\cdot;\bm{w})$ is a slowly varying function. A formal definition of $\Lambda(\cdot)$ is given in Appendix B.6. Note that the extended ADF corresponds with the copula exponent function introduced by Mackay and Jonathan, (2023) for data on uniform margins. The next proposition illustrates that, under mild conditions, the convergence in (6) is always achieved.
@@
Proof of Proposition 2.7 is provided in Appendix B.6. Note that while the assumptions of Proposition 2.7 may seem restrictive, Wadsworth and Tawn, (2013) demonstrate through rigorous theoretical treatment that model (2) captures the joint tail structure of $\bm{X}_{E}$ for a wide variety of theoretical examples. Therefore, it is reasonable to assume this framework can capture the tail structure in every orthant.
@@
In the definition of our model (6), we purposely exclude any angles in $\mathcal{S}^{d-1}$ that intersect the axes, as the model is not well defined there. Take, e.g., $\bm{w}:=(1,0,\ldots,0)^{T}$; it is not clear whether one should consider the probability $\Pr(X_{1}>u,X_{i}>0,i=2,\ldots,d)$ or $\Pr(X_{1}>u,X_{i}<0,i=2,\ldots,d)$ as $u\to\infty$ under the modelling framework. This results in discontinuities for the extended ADF at the axes; see Figure 2.
@@
Proof of Corollary 2.3 is provided in Appendix B.8, and visualised in Figure 2. The blue lines in each panel of Figure 2 denote the sets $\{\bm{w}/\Lambda(\bm{w}):\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}\}$, corresponding to the intersection points $\{\tilde{\mathfrak{r}}_{\bm{w}}\tilde{\mathcal{R}}_{\bm{w}}\cap\partial\mathcal{G},\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}$}. Furthermore, from the first and third panels of Figure 2, one can clearly observe the discontinuities of the extended ADF at the axes, as discussed in Remark 4.
@@ 3.2 Modelling assumptions for the conditional radii
Wadsworth and Campbell, (2024) demonstrate that, for large radial values $r$, we have that $f_{R\mid\bm{W}}(r\mid\bm{w})\propto r^{d-1}\exp\{-rg(\bm{w})\}$, where $f_{R\mid\bm{W}}(\cdot\mid\cdot)$ denotes the density function of $R\mid(\bm{W}=\bm{w})$. This implies that the upper tail of $R\mid(\bm{W}=\bm{w})$ follows a gamma kernel form. Note that this holds for $\bm{X}$ on both exponential and Laplace margins. To accommodate this form, Wadsworth and Campbell, (2024) propose the modelling assumption:
@@
where ‘truncGamma’ denotes a truncated Gamma distribution with shape and rate parameters $\alpha>0$ and $g(\bm{w})>0$, respectively, and $r_{\tau}(\bm{w})$ denotes the $\tau$-quantile of $R\mid(\bm{W}=\bm{w})$ for some $\tau\in(0,1)$ close to 1, that is, $\Pr\{R\leq r_{\tau}(\bm{w})\mid\bm{W}=\bm{w}\}=\tau$ for all $\bm{w}\in\mathcal{S}^{d-1}$. Under the distributional assumption of (7), we can use maximum likelihood techniques to estimate the gauge function $g(\cdot)$, which can be interpreted as a non-stationary rate parameter of the truncated gamma distribution.
@@ 3.3 Estimation of the gauge function using neural networks
To overcome the limited flexibility of existing geometric modelling approaches, we model the rescaled gauge function ($\tilde{g}$; see Equation (5)) using neural networks. Full inference requires the construction of two models: one for the radial threshold, $r_{\tau}(\bm{w})$ in (7), and one for an (unscaled) gauge function ${g}(\bm{w})$ satisfying ${g}(\bm{w})\geq||\bm{w}||_{\infty}$ for all $\bm{w}\in\mathcal{S}^{d-1}$ (see Proposition 2.2). With the latter, we can employ the transformation described in Equation (5), resulting in valid unit-level sets; see Corollary 2.2.
@@
where we have suppressed the dependency of $r_{\tau}(\cdot)$ on $\bm{\psi}$ in the notation. As $r_{\tau}(\bm{w})$ denotes the $\tau$-quantile of $R\mid(\bm{W}=\bm{w})$, the most appropriate choice of loss function is the tilted loss, given by $l(r,r_{\tau})=\rho_{\tau}(r-r_{\tau})$ for $\rho_{\tau}(z):=z(\tau-\mathbbm{1}\{z<0\})$ (Koenker et al.,, 2017). We can also define a suitable loss function and minimisation problem to train the deep representation of the rescaled gauge function, $\tilde{g}(\bm{w}).$ This model can be described as a conditional density network estimation (see, e.g., Rothfuss et al.,, 2019), where the loss function is the negative log-likelihood associated with the truncated gamma model defined in (7); note that this is dependent on both the exceedance threshold $r_{\tau}(\bm{w}_{j})$ and a scaling parameter, $\alpha>0$. To fit the gauge function model, we replace $\ell(r_{j},r_{\tau}(\bm{w}_{j}))$ in (8) with
@@
Full inference for our framework is performed in a two-stage fashion. We first train a model for $r_{\tau}(\bm{w})$ and derive its estimate, $\hat{r}_{\tau}(\bm{w})$. This estimate is then used in a subsequent training step for the rescaled gauge function $\tilde{g}(\cdot)$ and $\alpha$, by replacing $r_{\tau}(\bm{w})$ in (3.3) with its estimated counterpart. For details of the algorithms and regularisation techniques used for training of both models, as well as practical advice for pre-training (e.g., Goodfellow et al.,, 2016) of the models using initial estimates for $r_{\tau}(\cdot)$, see Appendix A.2.
@@ 3.4 Estimating the extended ADF
Estimates $\hat{\Lambda}(\bm{w}),\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A},$ can be used to obtain probability estimates for a wide variety of joint tail regions of $\bm{X}$. The estimation scheme outlined below avoids the need to sample from, or model, the distribution of the angles $\bm{W}$, as has been considered in other probability estimation schemes using the geometric representation (Papastathopoulos et al.,, 2024; Wadsworth and Campbell,, 2024). To begin, let $\bm{x}\in\mathbb{R}^{d}$ be such that $r:=||\bm{x}||$ is large and define the corresponding angle $\bm{w}:=\bm{x}/r$. Further define the structure variable $T_{\bm{w}}:=\min_{i=1,\dots,d}\{X_{i}/w_{i}\}$, and let $u$ denote a quantile of $T_{\bm{w}}$ satisfying $q:=\Pr(T_{\bm{w}}\leq u)<\Pr(T_{\bm{w}}\leq r)$, with $q$ close to 1 and $u>0$ large. Equation (6) implies that
@@
$=\Pr(T_{\bm{w}}>r\mid T_{\bm{w}}>u)\Pr(T_{\bm{w}}>u)\approx\exp\{-\hat{\Lambda}(\bm{w})(r-u)\}(1-q),$ (11)
where $\text{sgn}(\cdot)$ denotes the regular signum function. Figure 4 illustrates examples of joint tail regions in the two dimensional setting that can be estimated using the framework described in equation (6). Observe that these regions are more general than many inference procedures for multivariate extremes, which just focus on the joint survivor or cumulative distribution function (e.g., Ledford and Tawn,, 1996; Ramos and Ledford,, 2009).